SPECTROSCOPY AND STRUCTURE OF CHAINS TYPE C_{n}, C_{n}H and C_{n}Si

\underline{M.L. SENENT}, Departamento de Astrofísica Molecular e Infrarroja, Instituto de Estructura de la Materia, C.S.I.C., Madrid 28006, Spain.; M. HOCHLAF, Université Paris-Est, Laboratoire de Modélisation et Simulation Multi Echelle (FRE 3160 CNRS), Champs sur Marne, F-77454 Marnela-Vallée,Cedex 2, France.; H. MASSÓ, Departamento de Astrofísica Molecular e Infrarroja, Instituto de Estructura de la Materia, C.S.I.C., Madrid 28006, Spain.; N. INOSTROZA, Universidad Andrés Bello, Departamento de Química, Facultad de Ecología y Recursos Naturales, Santiago, Chile..

Structures, vertical excitation energies to the first electronic states and spectroscopic parameters for the ground electronic state are determined for several small C_{n}, C_{n}H and C_{n}Si astrophysically important molecular species ^{a}. The main aim of these investigations is to help the interpretation of several previous astrophysical observations ^{b} State-of-the art ab initio calculations are performed for this purpose. All species display isomerism although the inter-transformation processes, that involve various electronic states, are not well known. Anharmonic spectroscopic parameters are computed from full-dimensional Potential Energy Surfaces which are mapped close to their respective local minima. In light of these calculated properties (band positions, spectroscopic parameters, predictions of ro-vibronic effects, spin-spin constants,..) assignments for FIR astrophysical observed bands, are discussed. vspace1em

aH. Massó, M.L. Senent, P. Rosmus and M. Hochlaf, J.Chem.Phys. \underline{124}, 234304 (2006); M.L. Senent H. Massó and M. Hochlaf, ApJ, \underline{670}, 1510 (2007); H. Massó, V. Veryazov, P.A. Malmqvist, B.O. Roos and M.L. Senent, J.Chem.Phys. \underline{127}, 154318 (2007); M. Hochlaf, C. Nicolas and L. Poisson, J.Chem.Phys. \underline{127}, 014310 (2007) N. Inostroza, M. Hochlaf, M.L. Senent and J.R. Letelier (submitted 2008)

bJ. Goicoechea, J. Cernicharo, H. Massó and M.L. Senent, ApJ. \underline{609}, 225 (2004); J. Cernicharo, J.R. Goicochea and Y. Benilan, ApJ \underline{580}, L157, 2002.